OpenStudy (anonymous):

how do you find the maximum for r=3+4sin(theta)?

OpenStudy (angela210793):

by finding the 1st derivative r'=4cos(theta) r'=0 when cos(theta)=0 theta=+-pi/2 the max will be for theta=pi/2

OpenStudy (anonymous):

7

OpenStudy (anonymous):

find when sin theta is max..

OpenStudy (anonymous):

sin theta's max is 1 and minimum is -1 right? so do i multiply 4 to the one and then add 3?

OpenStudy (angela210793):

see the attachment :)

OpenStudy (anonymous):

wait...so is the answer pi/2 or 7..?

OpenStudy (angela210793):

theta must be pi/2 and for theta=pi/2 the max is r(pi/2)=3+4*1=7

OpenStudy (anonymous):

i think ans must be 7

OpenStudy (anonymous):

because the question is to find maxima for r

OpenStudy (angela210793):

u think theta=7? 7=r(pi/2)

OpenStudy (anonymous):

and theta=2npi+pi/2...

OpenStudy (angela210793):

yea ur right but i found wht should theta b then wht the max is

OpenStudy (anonymous):

i mean..r=7

OpenStudy (angela210793):

yea kk :)

OpenStudy (anonymous):

and theta has multiple solution if nothing has cited about the range of theta...

OpenStudy (angela210793):

well yea theta=2kpi+pi/2 but the angle would remain the same though

OpenStudy (anonymous):

so if i were to find the max of r=-3+2sin theta it would be 5? sorry im still a bit confused...

OpenStudy (anonymous):

where did the 2 come from in the -3+2=-1 ?

OpenStudy (angela210793):

no sorry theta=pi/6

OpenStudy (anonymous):

so the 2 is also the 2 in front of the sin in the original equation? i thought sin (pi/6) was 1/2 though..

OpenStudy (angela210793):

ur right!!!!!!!!!!!!!! I messed it up :(

OpenStudy (angela210793):

-3+1=-2 1 comes from 2*sin(pi/6)

OpenStudy (angela210793):

soooo soooooorrryyyyyyyyyy :(:(:( theta=pi/6 anyway :)

OpenStudy (anonymous):

o no it's fine! okay so the final answer would be -2? sorry i just want to make sure... so the first thing u do is find the max of the sin theta part. and then, you find out the value of the theta that'd satisfy the max. then, multiply that sin theta value to the number in front of the sin theta in the original. then, take the first number in the original (-3) and subtract the sin thetaxfront number to get the final answer...?

OpenStudy (anonymous):

sorry for asking so many questions! but i also want to know... in this case, r=4sin2(theta), max of sin2theta = 1 so sin2theta=1 theta= pi/4 sin pi/4 = sq2/2 max= 0+sq2/2 = sq2/2? i feel like i'm doing this wrong...

OpenStudy (angela210793):

sin2theta=1 2theta=pi/2 theta=pi/4 max=4*sin2(pi/4)=4sin(pi/2)=4

OpenStudy (anonymous):

thanks!

OpenStudy (angela210793):

uw